x 6.1 Class 8 Maths Question 1.
What will be the unit digit of the squares of the following numbers?
(i) 81
(ii) 272
(iii) 799
(iv) 3853
(v) 1234
(vi) 20387
(vii) 52698
(viii) 99880
(ix) 12796
(x) 55555

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
(i) Unit digit of 812 = 1 (ii) Unit digit of 2722 = 4 (iii) Unit digit of 7992 = 1 (iv) Unit digit of 38532 = 9 (v) Unit digit of 12342 = 6 (vi) Unit digit of 263872 = 9 (vii) Unit digit of 526982 = 4 (viii) Unit digit of 998802 = 0 (ix) Unit digit of 127962 = 6 (x) Unit digit of 555552 = 5